OpenStudy (maddy1251):

True or false: In(x+y)-In(x+y)=0. I think it is true. But I am not too sure.If it is true, can someone help me explain why?

OpenStudy (anonymous):

True :)

OpenStudy (chris_ec):

True

OpenStudy (chris_ec):

Anything subtracted from itself is zero.

OpenStudy (anonymous):

If you substitute x+y=u, you'll see that ln(u)-ln(u)=0 the same way that 1-1=0

OpenStudy (maddy1251):

Thank you guys!! If I could medal you all I would.

OpenStudy (chris_ec):

563blackghost, can you give CShrix a medal?